OpenStudy (anonymous):

x divided by -18? Does the x switch to a 1?

OpenStudy (anonymous):

@dan815 @jabez177 @Legends @rishavraj

rishavraj (rishavraj):

wht do u mean \[\frac{ x }{ -18 }\] hmmm so???

OpenStudy (anonymous):

Ya thats what I mean :)

rishavraj (rishavraj):

wht do u mean by x swithes to a ??

OpenStudy (anonymous):

x switched to 1 ? @rishavraj

rishavraj (rishavraj):

hmmmm x is a variable...... so it would be \[-\frac{ x }{ 18 }\]

OpenStudy (dan815):

okay listen up buttercup

OpenStudy (anonymous):

Me?

OpenStudy (dan815):

mhm :)

OpenStudy (dan815):

x divided by -18? Does the x switch to a 1? x/-18 x=a1?? whaT???!?!?!?

OpenStudy (anonymous):

Omg the sass in that

OpenStudy (dan815):

bahahah

OpenStudy (dan815):

hushhh u hussiiee

OpenStudy (dan815):

im still completely lost soo what is this a1?

OpenStudy (dan815):

ohh x switched to a 1 is x=1

OpenStudy (anonymous):

Yes but whatever I skipped it ;P

OpenStudy (dan815):

-1/18 = plug that into a calculator if u want decimal version of that

OpenStudy (dan815):

sigh do your work properly dummy

OpenStudy (anonymous):

:O Don't even get me started

OpenStudy (dan815):

hehe

OpenStudy (anonymous):

Remember the Squirrel dude?